Skip to main contentSkip to solution

A problem in Mathematics is given to two students X and Y whose chances of solving it are 13\frac{1}{3} and 14\frac{1}{4} respectively. The probability that only X solves the problem, is:

Solution

✅ Correct Option: 3

The probability that only X solves the problem means X successfully solves it and Y fails to solve it.

Given:

Probability that X solves the problem =13= \frac{1}{3}

Probability that Y solves the problem =14= \frac{1}{4}


The probability that Y does not solve the problem:

P(Y does not solve)=1−14P(\text{Y does not solve}) = 1 - \frac{1}{4}

P(Y does not solve)=34P(\text{Y does not solve}) = \frac{3}{4}


For only X to solve the problem, both events must occur: X solves AND Y does not solve.

Since these are independent events:

P(Only X solves)=P(X solves)×P(Y does not solve)P(\text{Only X solves}) = P(\text{X solves}) \times P(\text{Y does not solve})

P(Only X solves)=13×34P(\text{Only X solves}) = \frac{1}{3} \times \frac{3}{4}

P(Only X solves)=312P(\text{Only X solves}) = \frac{3}{12}

P(Only X solves)=14P(\text{Only X solves}) = \frac{1}{4}


Therefore, the probability that only X solves the problem is 14\frac{1}{4}.

Keyboard Shortcuts

  • Left arrow: Previous question
  • Right arrow: Next question
  • S key: Jump to solution
  • Q key: Jump to question